### v1

# A Multiobjective Mathematical Model for Optimal Irrigation Water Allocation

Abstract Water resource management is vital for sustainable irrigation and food security under growing climatic and economic pressures. In this study, we develop two single-objective optimization models for irrigation planning, one that maximizes net benefit and the other that minimizes environmental flow deficiency, and benchmark them against previous works. We then extend the analysis to a multiobjective linear programming formulation solved through different approaches to evaluate trade-offs. Numerical experiments on the Muhuri Irrigation Project reveal three outcomes: (i) a complete scenario view with profits ranging from $0.2\times 10^{9}$ to $1.497\times 10^{9}$ and environmental flow deficits between 0 and 1200 GL, where the 1200 GL represents the theoretical annual maximum under a 100 GL uniform monthly target; (ii) explicit trade-offs showing higher profits correspond to greater ecological shortfalls; and (iii) an integration-based approach producing nearly 1000 Pareto-optimal solutions within seconds, greatly outperforming earlier studies.
Keywords Irrigation Water Optimization, Environmental Flows, Crop Allocation, Pareto front, Muhuri Irrigation Project.
AMS subject classifications. 90B90, 90C29, 90C30, 92B05
1 Introduction
Although approximately 70 percent of the Earth’s surface is covered by water, only 2.5 percent constitutes fresh water, and less than 1 percent of that is accessible and usable for industrial, agricultural, and domestic purposes [18]. Specifically, only 0.0002 percent of water is available as surface water runoff, 0.62 percent exists as groundwater, and just 0.001 percent is stored in the atmosphere [16]. This severe imbalance between abundance and usability makes water scarcity one of the most critical global challenges of the present era. A sustainable future depends on effective resource management. Water resource management, therefore, requires a comprehensive and well-integrated approach. This approach should include the development of appropriate infrastructure, equitable distribution mechanisms, and the preservation of environmental flows necessary for rivers and ecosystems. Achieving these objectives is not possible without the support of robust institutions, advanced technologies, and sincere cooperation across all levels of society. Climate change, urban growth, and expanding agriculture influence the raising of water demand. Regarding this, managing water wisely has become more important than ever. Different regions face different challenges, so strategies must be tailored to local climates. For example, arid countries like Iran, Iraq, and Saudi Arabia must find ways to deal with severe water shortages and their heavy reliance on groundwater [21, 32]. In semi-arid regions like South Africa and Spain, irrigation systems need to be more flexible and drought-resistant [6, 10]. In contrast, tropical monsoon countries such as Bangladesh face the dual challenge of managing both excess rainfalls during monsoons and acute shortages in dry winter periods[28, 26].
A significant initiative aimed at improving water management in southeastern Bangladesh is the Muhuri Irrigation Project (MIP). This project was developed with a dam-regulator system to support dry-season irrigation and to prevent saline water from intruding inland from the Bay of Bengal. However, it has consistently underperformed due to system inefficiencies [25].Despite its strategic national importance, the MIP has received limited attention in academic research. Determining the optimal allocation of water for each crop, along with identifying suitable crop patterns under relevant constraints, remains a pressing research challenge.
Furthermore, in the previous study by Ullah and Nehring[28] reported a maximum Environmental Flow Deficiency (EFD) of 35 GL under a uniform monthly environmental flow target of 100 GL. Since EFD is defined as the nonnegative difference between the target environmental flow and the actual environmental flow, the theoretical maximum deficit occurs when no water is allocated to the environment. In such a case, the deficit would be equal to the full annual target of 100 GL × 12 months = 1,200 GL. Therefore, the reported value of 35 GL represents less than 3 percent of the possible range of deficits. This suggests that their model captured only a narrow portion of the feasible solution space and failed to approximate the full Pareto frontier of trade-offs between agricultural benefits and ecological requirements. As a result, the ecological shortfalls are understated, and the diversity of potential optimal solutions remains unexplored. The analysis is being extended to solve two single-objective problems individually: maximizing economic benefits and minimizing environmental flow deficits, which were not examined in previous studies. By employing alternative deterministic optimization approaches, including interior-point and sequential quadratic programming across different algorithms, we seek to generate accurate and reproducible outcomes that correct unrealistic allocations and reduce computational burden. These contributions not only enhance the validity of ecological and economic trade-offs but also extend the practical applicability of multiobjective water allocation models in resource-constrained irrigation projects.
The main objective of this research is to enhance the practical applicability and effectiveness of irrigation water allocation models in the Muhuri Irrigation Project. To achieve this, the following specific aims are proposed:
  • (i)
    Develop two single-objective mathematical models under a common set of constraints, including pumping balance, minimum area requirements, and total land availability. The objectives are formulated and solved individually: (a) maximization of economic benefit, and (b) minimization of environmental flow deficiency.
  • (ii)
    Compare established multiobjective methods with previous work on the Muhuri Irrigation Project using the same database[28].
  • (iii)
    Formulate single-objective mathematical models and obtain optimal solutions using interior-point and sequential quadratic programming across different algorithms.
  • (iv)
    Design efficient algorithms, evaluate their performance, and implement a structured solution procedure for the proposed models through computational experiments.
To address these pressing challenges, this study proposes a comprehensive optimization framework. The framework (i) evaluates and contrasts multiple optimization algorithms, and (ii) develops a robust decision-support system that closely reflects the real-world conditions faced in the field. The findings are expected to assist policymakers, irrigation planners, and water resource managers in making informed, data-driven, and climate-resilient decisions that enhance agricultural water efficiency and sustainability.
The rest of paper is organized as follows. Section 2 discusses the theoretical insights and managerial contributions of this study. Section 3 reviews previous works and highlights the problems encountered in the Muhuri Irrigation Project. Section 4 sets out the preliminary concepts used to build the proposed models, especially single objective and multiple objective models. Introduces two new mathematical formulations for constrained optimization problems, including environmental constraints. Section 5 presents numerical experiments and solutions for individual single objective models. Section 6 illustrates the Pareto front obtained from the solutions and compares the results with one another as well as with the previous study. Section 7 concludes with a summary of findings, limitations, and directions for future research.
2 Decision and Managerial Insights
The proposed multiobjective irrigation optimization model is well-suited for practical use in real-world irrigation planning the Muhuri Irrigation Project and similar initiatives, not just in Bangladesh but around the world. Combining seasonal water availability, the model offers a direct translation to field-level operations. It will assist the irrigation authorities to plan for different seasons, and allocate resources efficiently. The proposed irrigation optimization model serves a range of stakeholders. Farmers can use it to choose crop plans that balance income and water use, while irrigation managers and engineers can better schedule and regulate water flows. It also helps policymakers craft sustainable water-sharing strategies, and gives researchers a flexible tool to apply in other regions or for studying climate resilience. There are several practical benefits that can be gained by this model. First, the local community would be able to improve the crop productivity, and promote the equal sharing of the water resources. Economically, the model enhances profitability by making decision of minimum area of land and amount of water should be used for each crop, while concerning the ground water usage. Ecologically, it contributes to healthier river ecosystems by helping to reduce Environmental Flow Deficiency (EFD). For researchers, the model provides a replicable framework that can be adapted to other semi-arid regions facing similar challenges. This proposed model compared established multiobjective methods, e.g., gamultiobjective, pareto search, and scalarization techniques, with interior-point,sequential quadratic programming (SQP) which provides faster and more accurate optimization results. Unlike many traditional approaches, it takes into account several constraints in practical irrigation planning. By using both surface water and groundwater in a balanced way, the model helps reduce the cost of pumping water. It also helps protect crops from being damaged by overuse or underuse of water. In the end, this saves money by using land and farming inputs more wisely. The model is built to be practical and easy to use in real-life settings. It works well with common tools like AMPL, MATLAB, or Python. The required data are easy to collect—such as crop water needs, and seasonal water flow. Since it is designed in a simple scalarized form, it can also fit smoothly into the systems already used by water management authorities. It encourages fair sharing of water among farmers, which reduces conflicts and supports stronger, more resilient rural communities.
3 Literature Review
Efficient irrigation water management has become increasingly critical in balancing agricultural productivity with environmental sustainability. Modern research focuses not only on maximizing net benefits for farmers but also on ensuring long-term ecological health by minimizing environmental flow deficiency (EFD). This dual-objective perspective guides the evolution of optimization models in water resource planning. One of the earliest comprehensive efforts in multiobjective irrigation modeling was proposed by Xevi and Khan[31], who developed a model that simultaneously maximized farm income and minimized both groundwater extraction and variable costs. This dual-objective framework laid the groundwork for many subsequent models. Following a similar path, Lalehzari et al.[12] investigated optimal crop planning under water scarcity and emphasized how water use efficiency (WUE) could be improved through seasonal adjustments and pricing strategies. Their findings showed that vegetable crops consistently delivered the highest WUE. In the Bangladeshi context, Ullah and Nehring [28] applied a multiobjective approach to the Muhuri Irrigation Project (MIP), aiming to enhance economic returns while maintaining environmental sustainability. Although the model successfully produced 34 non-dominated solutions, it required extensive computation time. This gap reveals a need for more practically grounded frameworks in the context of Bangladeshi irrigation systems. In recent years, the inclusion of environmental objectives has become more prominent. Zeinali et al. [32] developed a model that balances surface and groundwater use across dry and wet periods, which led to reduced groundwater level decline—a key component of EFD minimization. Similarly, Hoque et al. [9] emphasized the importance of preserving environmental flows, particularly in dam-influenced areas, underscoring the need to harmonize dam operations with ecological flows. Adding practical constraints, Wardlaw and Bhaktikul [30] incorporated field-level soil moisture and canal capacity into their model, making it suitable for real-world implementation. Sadati et al. [24] examined both full and deficit irrigation strategies under four probabilistic weather conditions. They found that deficit irrigation consistently improved net farm income and total cropped area—contributing directly to benefit maximization while maintaining moderate water usage. Nguyen et al. [22] introduced a decision framework to reduce the computational search space in water allocation models. Their benchmark case study showed that this approach could improve solution efficiency, especially in complex, resource-intensive scenarios. Mohammadrezapour[19] evaluated different cropping options and reported that wheat provided the highest return-to-water ratio due to its low water requirement and high yield. At the same time, crops like sugar beet, which require high water input, saw significant decline in all modeled scenarios. His findings highlighted how crop selection impacts both water use and farm income. Similarly, Torabi [27] presented an integrated water operation model that improved water reliability, reduced groundwater drawdown, and minimized aquifer overuse, ensuring both short-term productivity and long-term resource sustainability. On a national scale, Musa [21] developed a model for Saudi Arabia that forecasts a shift from groundwater and surface water to treated wastewater and desalinated water between 2020 and 2050. The model projects water cost to rise from 13 Billion AUD to 22 Billion AUD, highlighting the importance of sustainable water reuse and quality management in long-term policy planning. In a more field-level approach, Bhatia and Rana [1] emphasized optimal crop allocation, resource efficiency, and sustainable practices like crop rotation to maintain soil health and ensure economic viability for farmers. While not entirely aligned with dual-objective modeling, some studies still offer useful insights. Ikudayisi [10], working in Baghdad’s Alrustumia district, analyzed the economic return from different types of treated wastewater. Though focused more on net benefit than environmental flow, the study demonstrated how wastewater reuse can support high-value crops. Marzban [15] proposed a flexible model for handling multiple goals in water-scarce regions but did not include explicit EFD constraints. Montazar [20] applied a model in Iran’s Qazvin Irrigation Command Area to evaluate cropping scenarios involving surface and groundwater use. While relevant in hydrological terms, the model lacked a strong focus on ecological sustainability. Some authors offer broader perspectives or critiques rather than direct model contributions. Lewis and Randall [13] highlighted limitations in earlier models, such as inadequate linkage between groundwater use and crop-specific water needs. Mainuddin [14] noted that Boro rice accounts for a major portion of groundwater use in Bangladesh, but did not incorporate optimization-based planning. Boah [2] pointed to the need for integrating water quality concerns into optimization frameworks. Kovalchuk [11] focused on managing river salinization through flushing techniques, important for pollution control, though outside the scope of crop-based irrigation allocation. Given the limitations in existing models, especially in the context of the Muhuri Irrigation Project, this research proposes efficient multiobjective methods such as ’gamultiobj’, ’paretosearch’, and ’scalarization’ approaches. The model integrates economic returns, seasonal water availability, and crop-specific water requirements. This approach seeks to deliver practical, efficient, and environmentally sound irrigation planning—bridging the gap between theoretical models and field-level applications in Bangladesh.
3.1 Problem Statement
The Muhuri Irrigation Project (MIP) is located in southeastern Bangladesh, covering parts of Feni, Chattogram, and Noakhali districts. The idea behind MIP took shape in the late 1980s, when the government looked for ways to increase crop output in the region. They used surface water from the Muhuri River, sending it through a network of canals and small field channels to reach the farms. But during the dry months, when supply drops, that’s when problems start. To deal with the shortage, many farmers start using groundwater as a backup option during those dry months. The project spans roughly 23076 hectares of irrigable area out of 40080 ha [29, Figure A.3], where farmers grow a variety of crops year-round, Boro rice, wheat, maize, and vegetables being the most common. Farmers prefer Rabi crops because these require less water than Kharif crops; however, in the dry winter season, production plummeted due to only 2 percent rainfall [28]. Seasonal variability in water supply and pressure to meet production targets frequently lead to inequities, with farmers at the tail end of canals receiving less water than those near the source. Ecologically, reduced river inflow during irrigation creates Environmental Flow Deficiency (EFD), threatening aquatic ecosystems, while excessive groundwater use depletes aquifers and raises energy costs. Given these realities, there is a pressing need for an updated planning model that integrates both economic objectives and ecological considerations. A robust decision-support system is essential to guide resource allocation and ensure the long-term sustainability of the Muhuri Irrigation Project.
4 Formulations of the Mathematical Models and Solution Approach
An irrigation planning authority must allocate land and water resources efficiently across multiple crops and time periods, ensuring economic viability and environmental sustainability. Surface water is delivered through canals, and additional water can be pumped. The system must meet staple crop area and production demands while minimizing environmental flow deficiency (EFD).
Decision variables:
$X_{c}$ Crop Area for each crop c in hectare haha
$Env.Flow_{m}$ Environmental Flow in month m in gigaliter GLGL
Input parameters:
$P_{c}$ Price of total crop production in Ton per hectare (T/ha)
$C_{w}$ Cost of Surface Water conveyance (AUD/GL)
$K_{cm}$ Crop coefficient in the different growth stages of crops c in month m
$ET_{m}$ Evapotranspiration($GL/ha$) in each month m
$R_{m}$ Rainfall($GL/ha$) in each month m
$Allocation_{m}$ Surface Water($GL$) in each month m
$P_{m}$ Ground or Pumping Water($GL$) in each month m
$C_{p}$ Cost of pumping Groundwater(AUD/GL)
$Vcost_{c}$ Variable cost for each crop c per hectare (AUD/ha)
$Inflow_{m}$ Total runoff or Available River water ($GL$)in each month
$Tar.Env.Flow_{m}$ The minimum required Environmental flow($GL$) in each month
$T_{\text{pump}}$ Total Pumping Water $(GL)$
$T_{\text{area}}$ Total Area $(ha)$
$min_{area}$ Minimum Area $(ha)$
Decision variables dependent parameters:
$Allocation_{m}$ Allocated Surface water, where $Allocation_{m}={Inflow}_{m}-{Env.Flow}_{m}$
$P_{m}$ Pumped Ground Water, where $P_{m}=\sum_{c}(K_{cm}\cdot ET_{m}-R_{m})X_{c}-Allocation_{m}$
The objective functions can be formulated as follows:
$NB_{\text{max}}=\sum_{c}P_{c}X_{c}-C_{w}\cdot\sum_{m}\left(\sum_{c}(K_{cm}\cdot ET_{m}-R_{m})X_{c}-P_{m}\right)\\ -C_{p}\cdot\sum_{m}P_{m}-\sum_{c}Vcost_{c}\cdot X_{c}$ (1)
The second objective, $f_{2}$:
$EFD_{\text{min}}=\sum_{m}\max\left(\text{Tar.Env.Flow}_{m}-\text{Env.Flow}_{m},\ 0\right)$ (2)
Model 1: When Net Benefit is maximized
$\begin{array}[]{rl}\displaystyle\max_{(X_{c},Env.Flow_{m})\in X}&\ f_{1}(X_{c},Env.Flow_{m})\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (3)
where the feasible set $X$ is defined as
$\hskip-14.22636ptX:=\left\{(X_{c},Env.Flow_{m})\in\mathbb{R}^{c}\times\mathbb{R}^{m}\mid\sum_{m}P_{m}\leq T_{pump},\;\;min_{area}\leq X_{c},\;\;\sum_{c}X_{c}\leq T_{area}\right\}.$
The model seeks to maximize net benefit, with $X_{c}$ and $Env.Flow_{m}$ as the decision variables and parameters, including $P_{c}$, $C_{w}$, $K_{c,m}$, $ET_{m}$, $R_{m}$, $Allocation_{m}$, $P_{m}$, $C_{p}$, and $Vcost_{c}$. In this setup, maximum net benefit is achieved by sacrificing environmental performance, illustrating the trade-off between economic gain and ecological sustainability.
Model 2: When Environmental Flow Deficiency is minimized
$\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ f_{2}(X_{c},Env.Flow_{m})\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (4)
Hence, our consideration is to minimize the Environmental Flow Deficiency (EFD). Specifically, if the target environmental flow is set equal to or lower than the actual environmental flow, the EFD becomes zero. This restriction ensures that minimizing EFD is achieved, though it results in the minimum attainable net benefit.
Model 3: Multiobjective - Net Benefit versus Environmental Flow Deficiency
$\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\displaystyle\ \left\{-f_{1},f_{2}\right\}\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (5)
This multiobjective programming model simultaneously minimizes two conflicting goals: the negative Net Benefit function $f_{1}(X_{c},Env.Flow_{m})$ and the environmental flow deficiency function $f_{2}(X_{c},Env.Flow_{m})$. The constraints ensure feasible pumping limits, enforce both minimum and maximum land-use requirements, and guarantee that crop area and environmental flow allocations remain non-negative. In essence, the formulation captures the trade-off between maximizing agricultural returns and preserving ecological water requirements.
We develop code within the MATLAB environment and utilise MATLAB solvers to address the models. A comprehensive solution procedure and algorithm for solving Models 1–3 are detailed in Appendices 9.1 and 9.2.
5 Numerical Experiments, Results and Sensitivity Analysis
This section presents a numerical study that shows the optimal water allocation for irrigation can be achieved under a large number of constraints. The solution procedure given in (9.1) is used to maximize profit for Models (3) and minimize environmental deficiency of Model (4). MATLAB solvers are used, including ’fmincon’ with the interior-point method and the sequential quadratic programming (SQP) algorithm to solve the problems. For the multiobjective problem, we employ the MATLAB genetic algorithm and the Pareto search approach, along with the scalarization method described in [3]. These solvers are tested for their ability to handle large-scale optimization and the challenges of approximating the Pareto front. Among the tested options, we report the computational time taken and the approximation of the solution of the solvers used. All computations are carried out on an HP Evo laptop with $16$ GB RAM and a Core $i7$ processor running at $4.6$ GHz.
Case Study 5.1
This numerical example is based on the case study by Ullah and Nehring [28]. The study considers a regional water board managing ten main crops over 12 months. The total cultivable area is 23,076 hectares, with a maximum annual water pumping capacity of 50 GL and a minimum crop demand for each 1000 hectares. The costs for pumped water and surface water are 100,000 and 26,000 AUDAUD perper GLGL, respectively. The target environmental flow is set at 100 GL per month. Additional parameter data, including rainfall, evapotranspiration, crop coefficients, crop production, total crop income, and variable costs, are provided in [28, 29]. We are examining how to allocate crops across the available area while considering whether the environmental flow meets its target environmental flow requirement. We analyze cases both with and without satisfying this constraint to determine the highest possible profit. This analysis provides insights that can later be used to explore the trade-off between maximising profit and minimising environmental deficiency.
Table 1: Optimal Allocation of Crops and Environmental Flow for Model (3).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 2076 1000 1000 5000 5000 5000
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 0 0 0 0 0 0 0 0 0 0 0 0
We solved Model 3 and obtained the crop and environmental flow allocations shown in Table 1. In this case, the maximum profit achieved is $1.4968\times 10^{9}$. At this point, the maximum deficiency in environmental flow is $1,200$ GL.
Now we consider that the environment flow meets the target flow requirement of the environment in Model (3), that is,
$Tar.Env.Flow_{m}\leq Env.Flow_{m},$ (6)
Table 2: Optimal crop and environmental flow allocation for Model (3) under constraint (6).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 2076 1000 1000 5000 5000 5000
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 100 100 100 100 100 100 100 100 100 100 100 100
We solved Model 3 using constraint (6) and obtained the crop and environmental flow allocations presented in Table 2. In this scenario, the maximum profit is $1.4080\times 10^{9}$. Here, the environmental flow deficiency is 0 GL, since constraint (6) does not allow for any shortfall.
We next minimize Model 4 with the same parameter settings as in Case Study 5.1, considering both the inclusion and exclusion of constraint (6). The optimal crop and environmental flow allocations, shown in Table 3, are identical in both situations. In this case, the total environmental flow deficiency is zero, and the profit obtained is $0.2742\times 10^{9}$. Note: The computational time taken to solve all three cases is approximately 8–10 seconds.
Table 3: Optimal crop and environmental flow allocation for Model (4).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 1000 1000 1000 1000 1000 1000
Env. Jan Feb Mar April May June July Aug Sep Oct Nov Dec
Flow 200 200 200 200 200 200 200 200 200 200 200 200
Observations from the Results: The three scenarios deliver different outcomes because each goal controls how water is shared between crops and the river:
  • Profit only: When the model focuses only on profit, nearly all available water is directed to the most valuable crops. This generates the highest income of $\mathdollar 1.4968\times 10^{9}$, but the river receives too little water, creating a total environmental flow shortage of $1,200$ GL.
  • Profit with environmental target: When the model must also meet the river’s target flow, it reserves enough water for the environment. With less surface water left for irrigation, more pumping is needed, which is expensive. As a result, profit drops moderately to $\mathdollar 1.4080\times 10^{9}$, but the environmental shortage is reduced to zero.
  • Minimum environmental shortage: When protecting the environment is the top priority, the model sends a steady 200 GL each month to the river before allocating water to crops. Because of this and limits on pumping, not all land can be irrigated, and profit falls sharply to $\mathdollar 0.2742\times 10^{9}$. The river’s flow target, however, is fully met.
Overall, these results highlight a clear trade-off: directing more water to crops increases income but harms ecological health, while stronger environmental protection lowers profit.
6 Multiobjective Optimization Problems
Multiobjective optimization deals with finding the best balance between two or more objectives that often compete with each other. Because these goals conflict, there is rarely a single solution that is best for every objective at once. Instead, the aim is to identify a group of equally valid trade-off solutions called the Pareto front, where improving one objective would worsen at least one other. A common way to address such problems is scalarization, which reformulates a multiobjective problem as a single-objective one so that standard optimization techniques can be applied. In this work, we use the weighted-constraint method [3]. Here, the objective functions are expressed with adjustable weights: one weighted objective is optimized, while the remaining objectives are treated as constraints whose bounds depend on the chosen weights. By varying these weights evenly, we can approximate the complete Pareto-optimal set. We chose the weighted-constraint approach for its straightforward implementation and its ability to generate accurate Pareto solutions (more scalarization approaches in [23][4]). The detailed procedure for the two-objective case is described below.
The weighted-constraint approach ($P_{w}^{k}$) : For a positive weights we define
$\hskip 56.9055ptW:=\left\{w\in\mathbb{R}^{2}\mid w_{1},w_{2}>0,\sum_{i=1}^{2}w_{i}=1\right\}.$
Let $w\in W$, $f_{1}$ in (1), and $f_{2}$ in (2), and we therefore define two scalar problems as
$\mbox{($P_{w}^{1}$)}\ \left\{\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ \ w_{1}f_{1}(X_{c},Env.Flow_{m}),\\[11.38109pt] \mbox{subject to}&\ \ w_{2}f_{2}(X_{c},Env.Flow_{m})\leq w_{1}f_{1}(X_{c},Env.Flow_{m}),\\ &\ \ \displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}\right$. (7)
and
$\mbox{($P_{w}^{2}$)}\ \left\{\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ \ w_{2}f_{2}(X_{c},Env.Flow_{m}),\\[11.38109pt] \mbox{subject to}&\ \ w_{1}f_{1}(X_{c},Env.Flow_{m})\leq w_{2}f_{2}(X_{c},Env.Flow_{m}),\\ &\ \ \displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}\right$. (8)
In Section 5, net profit and environmental deficiency are optimized separately, and the results highlight the trade-off between these two objectives. In Case Study 6.1, we solve the two-objective optimisation problem defined in (5).
Case Study 6.1
In this analysis, we use the input parameters from Case Study 5.1 to solve the multiobjective optimization problem defined in Model 3. Because it is a large-scale problem, different approaches are applied, including the ’gamultiobj’ and ’paretosearch’ solvers in the MATLAB environment, while the weighted-constraint method is employed to approximate the Pareto front.
We solve the problem described in (5) using the ‘gamultiobj‘ solver, increasing the population size up to $500$. The lower bounds are set as the minimum area for $X_{c}$ and zero for environmental flow, while the upper bounds are set at $5,000$ ha for $X_{c}$ and $300$ GL for environmental flow. The Pareto front obtained with this algorithm is shown in Figure 1(a), and the computation time along with the total number of Pareto points is reported in Table 4.
We now solve the problem defined in (5) using the ‘paretosearch‘ solver, with the population size increased to 500. The same lower and upper bounds are considered for $X_{c}$ and environmental flow values. The Pareto front approximated by this algorithm is shown in Figure 1(b), while the computation time and the total number of Pareto points are reported in Table 4.
Finally, we use the weighted-constraint scalarization approach. This method requires a sequence of steps detailed in Appendix 9.2. With weight normalisation as defined in $W$, the procedure involves solving subproblems (7) and (8) and comparing their solutions, discarding any non-dominated points. To solve the subproblems, we employ the fmincon solver with the ’Interior-Point’, ’SQP’, and ’SQP-Legacy’ algorithms, using the same lower and upper bounds as in the previous two cases. Following the algorithm outlined in Appendix 9.2, we then solve the problem defined in (5) to approximate the Pareto points, which are illustrated in Figure 1(c). A numerical summary is provided in Table 4.
Table 4: Numerical performance of ’gamultiobj’, ’paretosearch’ and $P_{w}^{k}$.
Methods Number of Population Number of Time Pareto Non-Pareto
grid points size subproblems in seconds points points
gamultiobj 500 15.72 345 Nil
paretosearch 500 16.86 500 Nil
$P_{w}^{k}$ 500 1000 10.12 989 11
Comparison Analysis: We observe that the ’gamultiobj’ algorithm generates $345$ Pareto points in $15.72$ seconds under the same parameter settings. From Figure 1(a), the highest environmental deficiency level is around $130$ GL when profit is $1.91\times 10^{9}$, and the deficiency drops to zero when profit is between $(1.180\;\mbox{and}\;1.181)\times 10^{9}$. The ’paretosearch’ algorithm takes slightly more time but produces more Pareto points $500$ in total. In generating the Pareto front, what matters more than the number of points is how much of the Pareto front each algorithm can cover. For paretosearch, the highest deficiency is $1200$ GL at a profit of $1.49\times 10^{9}$, with deficiency reaching zero when profit is around $1.37\times 10^{9}$ (see, Figure 1(b)). Using the weighted-constraint scalarization approach, we find that the highest deficiency and corresponding profit are the same as ’paretosearch’. However, the deficiency starts to appear when the profit is $1.4\times 10^{9}$ and decreases to $0.2\times 10^{9}$ (see, Figure 1(c)).
To evaluate the true Pareto front, we refer to Case Study 5.1 results, where individual objectives $f_{1}$ and $f_{2}$ are optimised. The results show that the highest deficiency of $1200$ GL occurs at a profit of $1.49\times 10^{9}$, consistent with Figures 1(b) and (c) obtained for ’paretosearch’ and the ’scalarization’ approach, respectively. Moreover, maximising $f_{1}$ under target environmental flow constraints (6) in Case Study 5.1 results in zero deficiency when profit is $1.4\times 10^{9}$, which matches only the scalarization results in Figures 1(c). This indicates that ’paretosearch’ and ’gamultiobj’ fail to approximate the lower part of the Pareto front. Therefore, considering computation time, the number of generated points, and the accuracy of the true Pareto front, the scalarization approach performs better than the other methods.
Comparison analysis with the previous study [28]: The findings of the present study highlight several critical improvements over the earlier work of Ullah and Nehring [28]. A key point of divergence lies in the treatment of Environmental Flow Deficiency (EFD). Ullah and Nehring reported a maximum EFD of only 35 GL under a fixed monthly environmental flow target of $100$ GL. Given that EFD is defined as the shortfall between the prescribed environmental flow and the actual allocation, the theoretical maximum deficit for a uniform 100 GL monthly target is $1,200$ GL per year (see, obtained results in Table 1). This indicates that their model captured only a narrow portion of the feasible solution space. Consequently, the ecological shortfalls are understated, and the trade-offs between agricultural benefits and ecological requirements are not fully explored. In contrast, our scalarization-based approach succeeds in covering the entire feasible frontier, generating Pareto solutions that extend up to the theoretical maximum EFD. This broader coverage, as illustrated in Figure 1(c) (Pareto fronts generated by the Weighted-Constraint approach, $P_{w}^{k}$), provides a far more comprehensive representation of the trade-off surface compared to [28, Figure 1 in Page 10], where the solution set remains confined to a very limited region.
The two studies also differ markedly in terms of computational performance and solution diversity. Ullah and Nehring, using the NSGA-II evolutionary algorithm, reported only $34$ Pareto-optimal solutions, obtained at the expense of approximately $706.59$ minutes of computational time. In sharp contrast, the scalarization technique produces $989$ Pareto solutions in just $10.12$ seconds. This demonstrates not only substantial computational efficiency but also a significant improvement in the diversity of solutions available to decision-makers. The larger and more representative solution set provided by our approach ensures a more reliable understanding of trade-offs and expands the range of strategies that can be considered in irrigation planning.
Together, these results establish that the scalarization framework achieves superior performance in both accuracy and efficiency. By covering the full spectrum of environmental deficits and delivering a wide set of optimal solutions with minimal computational burden, our study offers a more rigorous and practical foundation for decision-making in the management of the Muhuri Irrigation Project.
(a) Pareto front approximated with ’gamultiobj’.
(b) Pareto front approximated with ’paretosearch’.
(c) Pareto front approximated with ($P_{w}^{k}$).
Figure 1: Pareto fronts generated by the ’gamultiobj’, ’paretosearch’, and ’weighted-constraint’ ($P_{w}^{k}$) methods.
7 Conclusion
This study contributes to the growing efforts in sustainable water management by introducing a practical optimization framework designed to strike a balance between economic gain and environmental responsibility. By using three widely used techniques, gamultiobjectivite,Paretosearch and scalarization, the model provides an accessible and effective tool that can support planners in making more informed and thoughtful decisions when managing scarce water resources. The findings suggest that smart allocation strategies can lead to noticeably higher economic benefits while also safeguarding environmental flows. The results show that scalarization, in particular, is well-suited to handle trade-offs between conflicting goals. The model can also successfully consider real-world operational constraints, making it relevant and applicable in practice. Moreover, since the analysis is focused on a single irrigation project in a tropical monsoon area, the findings are also applicable in arid and semi-arid areas.
The study has some limitations on the type of data used. It follows a deterministic model where data are certain and previously known. It doesn’t fully reflect uncertainties or covers Stochastic or fuzzy aspects like changing climate conditions, fluctuations in crop yields, or shifting policy and social dynamics.
There is a scope to enhance the model by including different area boundaries for each crop on its demand, canal capacity constraints and crop scheduling based on land and crop types, as well as by adding deep percolation, soil moisture, dam storage, overflow, and reservoir evaporation constraints in future work. In addition, connecting the model with socio-economic and policy frameworks could help to better understand its broader impact and make it even more useful for decision-makers at both local and national levels.
Acknowledgment: We acknowledge Chatgpt(5) AI tools, which are used to rephrase, edit and polish the author’s written text for spelling, grammar, or general style.
8 Funding and/or Conflicts of interests/Competing interests
The authors declare that no funds, grants, or other support was received during the preparation of this manuscript. The authors have no relevant financial or non-financial interests to disclose.