### v2

# A Multiobjective Mathematical Model for Optimal Irrigation Water Allocation

Abstract Sustainable irrigation and food security increasingly depend on efficient water resource management in the face of growing climatic and economic constraints. In this study, we develop two single objective optimization models for irrigation planning, one that maximizes net benefit and the other that minimizes environmental flow deficiency, and compare their performance with established models reported in previous studies. We then extend the analysis to a multiobjective programming formulation solved through scalarization and genetic approaches to evaluate trade-offs. Numerical experiments on the Muhuri Irrigation Project reveal three outcomes: (i) a complete scenario view with profits ranging from $\mathdollar 0.2\times 10^{9}$ to $\mathdollar 1.497\times 10^{9}$ and environmental flow deficits between 0 and 1200 GL, where the 1200 GL represents the theoretical annual maximum under a 100 GL uniform monthly target; (ii) explicit trade-offs showing higher profits correspond to greater ecological shortfalls; and (iii) an integration based approach producing nearly 1000 Pareto optimal solutions within seconds, greatly outperforming earlier studies.
Keywords Irrigation Water Allocation, Environmental Flows, Crop Allocation, Multiobjective Optimization, Muhuri Irrigation Project.
AMS subject classifications. 90B90, 90C29, 90C30, 92B05
1 Introduction
Irrigated agriculture remains one of the largest consumers of freshwater worldwide, particularly in regions where seasonal variability and competing demands place sustained pressure on limited water resources. Although approximately 70% of the Earth’s surface is covered by water, only 2.5% constitutes fresh water, and less than 1% of that amount is accessible for industrial, agricultural, and domestic use [17]. Of this total, only 0.0002% is available as surface water runoff, while 0.62% exists as groundwater, and just 0.001% is stored in the atmosphere [15]. This imbalance between water availability and practical usability has made effective irrigation water allocation a critical challenge for ensuring food security, sustaining rural livelihoods, and protecting aquatic ecosystems. In many agricultural systems, irrigation planning must reconcile competing objectives: maximizing crop production while maintaining sufficient river flows to sustain ecological processes. These challenges are intensified by climate change, population growth, and expanding cultivation, which together increase water demand while reducing the reliability of surface water supplies. As a result, irrigation systems increasingly rely on conjunctive use strategies, where surface water is supplemented by groundwater during dry seasons. Under such conditions, allocation decisions are no longer purely operational but become strategic planning problems with long term economic, environmental, and social consequences.
Environmental flow, or e-flow, refers to the quantity, timing, and quality of water required to sustain the ecological integrity and functioning of a river system. Rather than being a fixed value, environmental flow varies seasonally in accordance with the river’s natural flow regime, supporting key processes such as fish migration, floodplain inundation, wetland regeneration, and groundwater recharge. Excessive water abstraction for agriculture or industry disrupts this natural rhythm, leading to ecosystem degradation, reduced recharge, and long term environmental and social costs. Maintaining minimum river flows is therefore not only an ecological requirement but also an economic and societal necessity. Any shortfall below the required environmental flow level is defined as Environmental Flow Deficiency (EFD), the magnitude of which depends critically on how realistically and flexibly environmental flow targets are specified within planning frameworks.
These challenges are particularly evident in monsoon dominated agricultural systems. Countries such as Bangladesh experience excessive rainfall during monsoon months alongside acute water shortages in the dry winter season. Irrigation systems in these regions must therefore manage both surplus and scarcity while supporting crop production across highly contrasting hydrological conditions. Allocation decisions directly influence cropping patterns, cultivated area, seasonal yields, and farmer livelihoods, making robust planning tools essential for agricultural sustainability [24, 26].
A major initiative to improve water management in southeastern Bangladesh is the Muhuri Irrigation Project (MIP). The project was developed with a dam regulator system to support dry season irrigation while preventing the intrusion of saline water from the Bay of Bengal into inland areas. However, it has consistently underperformed due to system inefficiencies[26]. Effective irrigation systems depend on sound water resource management, and in this context, the MIP was designed to utilize river water for agricultural production through planned and controlled distribution. Over the past several decades, the Bangladesh Water Development Board (BWDB) has implemented numerous water resource development and management programs, among which the Muhuri Irrigation Project represents a targeted effort to enhance agricultural productivity through improved irrigation management. Despite its strategic importance at the national level, the MIP has received limited attention in academic research. Consequently, determining the optimal allocation of water among crops and identifying suitable cropping patterns under relevant constraints remains an important and unresolved research challenge.
A key limitation of existing allocation studies for the MIP lies in their restricted representation of decision space. Previous work has evaluated irrigation performance primarily within a narrow range of environmental flow deficits. For example, Ullah and Nehring [26] reported a maximum Environmental Flow Deficiency (EFD) of 35 GL under a uniform monthly environmental flow target of 100 GL. Given that EFD is defined as the non negative shortfall between target and actual environmental flow, the theoretical maximum deficit corresponding to zero environmental release would equal the full annual target of 1,200 GL. The reported value therefore represents less than 3% of the feasible deficit range, indicating that only a limited portion of the potential economic–environmental trade-off space has been explored.
This narrow focus constrains decision relevance. When irrigation systems are pushed beyond moderate operating conditions such as during prolonged dry seasons or under policy driven environmental constraints planners lack guidance on how economic benefits evolve across the full range of possible environmental shortfalls. As a result, allocation decisions risk understating ecological stress or overestimating economic resilience, limiting their usefulness for real world agricultural planning.
To address this limitation, this study aims to improve the computational representation of trade-offs between economic and environmental objectives in irrigation allocation models, using the Muhuri Irrigation Project as a benchmark case study. Specifically, the study expands the decision space by explicitly characterizing trade-offs between net economic benefit and environmental flow deficiency across a substantially wider allocation range. By developing and comparing complementary optimization formulations under a consistent set of operational constraints, the study enables systematic evaluation of economic–environmental trade-offs under stylized seasonal water scarcity conditions.
The main objective of this research is to enhance the practical applicability and effectiveness of irrigation water allocation models in the Muhuri Irrigation Project. To achieve this, the following specific aims are proposed:
  • (i)
    Develop two single objective mathematical models under a common set of constraints, including pumping balance, minimum area requirements, and total land in the MIP. The objectives are formulated and solved individually: (a) maximization of economic benefit, and (b) minimization of environmental flow deficiency.
  • (ii)
    Compare established multiobjective methods with previous work on the Muhuri Irrigation Project using the same database[26].
  • (iii)
    Formulate single objective mathematical models and obtain optimal solutions using interior point and sequential quadratic programming across different algorithms.
  • (iv)
    Design efficient algorithms, evaluate their performance, and implement a structured solution procedure for the proposed models through computational experiments.
The rest of paper is organized as follows. Section 2 discusses the theoretical insights and managerial contributions of this study. Section 3 reviews previous works and highlights the problems encountered in the Muhuri Irrigation Project. Section 5 sets out the preliminary concepts used to build the proposed models, especially single objective and multiple objective models. Introduces two new mathematical formulations for constrained optimization problems, including environmental constraints. Section 6 presents numerical experiments and solutions for individual single objective models. Section 7 illustrates the Pareto front obtained from the solutions and compares the results with one another as well as with the previous study. Section 8 concludes with a summary of findings, limitations, and directions for future research.
2 Decision and Managerial Insights
This study presents three mathematical optimization models for irrigation water allocation in an irrigation system that relies primarily on surface water while supplementing supply with groundwater during periods of seasonal scarcity. The models are designed to represent alternative planning priorities and to clarify trade-offs between economic performance and environmental flow protection under constrained water availability. Model 1 emphasizes the maximization of net economic benefit, highlighting allocation strategies that favor agricultural production but may increase environmental flow deficit. Model 2 prioritizes the reduction of environmental flow deficiency, supporting ecological sustainability at the expense of lower economic returns. Model 3 integrates both objectives within a multiobjective optimization framework, allowing economic and environmental outcomes to be evaluated simultaneously and supporting balanced allocation decisions.
A key contribution of this study is the expansion of the decision space beyond the limited operating ranges commonly examined in earlier planning efforts. Previous analysis on MIP largely focused on moderate levels of environmental flow deficit, restricting managerial insight to a narrow set of near optimal outcomes. In contrast, the proposed framework illustrates how net economic benefits evolve across a much wider range of environmental flow deficits within a short evaluation horizon. This extended representation enables farmers and local authorities to assess not only favorable allocation options but also the economic implications of operating under increasingly stressed conditions, which are often encountered during dry seasons.
The proposed models further enhance decision support by treating both crop land allocation and environmental flow release as endogenous decision variables rather than fixed parameters. This formulation allows the optimization process to determine optimal crop area distribution and environmental flow allocation simultaneously, subject to shared operational constraints such as total cultivable land, pumping balance, and minimum area requirements. As a result, the interaction between agricultural production and ecological protection emerges directly from the model structure rather than being imposed through predefined allocation rules.
The multiobjective formulation explicitly characterizes trade-offs between net economic benefit and environmental flow deficiency by generating a Pareto set of feasible solutions. Instead of relying on ad hoc weighting schemes, this approach allows decision makers to evaluate multiple compromise strategies and select allocation policies that align with shifting priorities across seasons, hydrological conditions, or policy objectives. By making these trade-offs transparent, the framework supports more defensible and adaptive planning decisions in contexts where economic and environmental goals compete.
From a managerial perspective, the proposed models function as a practical decision support tool for irrigation planners, water resource managers, and policymakers. The framework enables users to compare allocation strategies across the full operating range, identify zones of diminishing economic returns, and understand the consequences of pushing the system toward higher levels of environmental stress. This reduces reliance on trial and error decision making and supports quicker, more informed responses under water scarcity and institutional pressure.
The analysis is conducted entirely using secondary data previously published for the Muhuri Irrigation Project, ensuring methodological consistency and direct comparability with earlier studies. All hydrological inputs, crop parameters, and environmental flow requirements are adopted without modification. Accordingly, the contribution of this study lies in improving solution robustness, efficiency, and trade-off representation through alternative optimization formulations rather than in site specific data recalibration or operational redesign.
The scope of the study is clearly defined. While the models are calibrated for the Muhuri Irrigation Project, the underlying framework is applicable to other regulated irrigation systems that rely primarily on surface water while supplementing supply with groundwater during dry periods. At the same time, the conclusions remain grounded in local hydrological, institutional, and socioeconomic conditions, ensuring relevance without overgeneralization.
3 Literature Review
Efficient irrigation water management has become increasingly important as agricultural systems face growing pressure to balance economic productivity with environmental sustainability. Contemporary irrigation planning therefore extends beyond profit maximization to include ecological objectives, particularly the preservation of riverine health through environmental flow protection. A key metric used to represent this ecological requirement is Environmental Flow Deficiency (EFD), defined as the non negative shortfall between a target environmental flow and the actual flow released to the river. When actual releases fall below the ecological target, the deficit contributes to EFD, while exceedances do not incur penalties. This formulation provides a transparent and quantifiable way to incorporate ecological constraints into irrigation water allocation models.
One of the earliest comprehensive multiobjective optimization frameworks for irrigation planning was proposed by Xevi and Khan [29], who developed a model that simultaneously maximized farm income while minimizing groundwater extraction and variable irrigation costs. Their work established the foundation for integrating economic and hydrological objectives within a single optimization framework. Building on this approach, Lalehzari et al. [11] investigated optimal crop planning under water scarcity and demonstrated how water use efficiency could be improved through seasonal adjustments and pricing mechanisms. Their results showed that vegetable crops consistently achieved higher water use efficiency, reinforcing the importance of crop selection in water allocation decisions. These early studies collectively demonstrated the potential of multiobjective optimization in irrigation planning, although explicit ecological flow targets were not yet a central modeling focus.
Subsequent research increasingly emphasized environmental considerations in irrigation modeling. Zeinali et al. [30] developed a framework that balanced surface and groundwater use across wet and dry seasons, demonstrating that coordinated use could significantly reduce groundwater depletion, a key contributor to environmental flow preservation. Hoque et al. [8] further highlighted the importance of maintaining environmental flows in dam regulated systems, underscoring the need to harmonize dam operations with downstream ecological requirements.
At the operational level, Wardlaw and Bhaktikul [28] incorporated field scale constraints such as soil moisture dynamics and canal capacity, thereby improving the realism and applicability of irrigation optimization models. Sadati et al. [23] examined full and deficit irrigation strategies under probabilistic weather conditions and showed that deficit irrigation could improve net farm income while maintaining moderate water use. Nguyen et al. [21] contributed methodologically by proposing a framework to reduce the computational search space in large scale allocation problems, improving solution efficiency in complex systems.
Several studies examined the influence of crop selection on both economic returns and water use. Mohammadrezapour [18] reported that wheat achieved the highest return to water ratio due to its relatively low water requirement and high yield, while water intensive crops such as sugar beet performed poorly under scarcity. Torabi [25] proposed an integrated operational model that improved water reliability, reduced groundwater drawdown, and minimized aquifer overuse, emphasizing the importance of coordinated surface and groundwater management.
At a broader policy scale, Musa [20] developed a national water planning model for Saudi Arabia, projecting a long term shift toward treated wastewater and desalinated water sources. Although contextually different, the study highlighted the economic consequences of unsustainable water use. Complementing this perspective, Bhatia and Rana [1] emphasized optimal crop allocation and sustainable farming practices such as crop rotation to maintain soil health and farm viability.
Several studies provide important conceptual foundations but do not explicitly incorporate environmental flow deficiency. Ikudayisi [9] analyzed economic returns from treated wastewater reuse in Iraq, demonstrating benefits for high value crops but without explicit ecological flow constraints. Montazar [19] evaluated cropping scenarios using both surface and groundwater in Iran’s Qazvin Irrigation Command Area, though ecological sustainability was not a primary objective. Marzban [14] proposed a flexible multi goal optimization framework for water scarce regions but did not explicitly model EFD.
Lewis and Randall [12] made a key methodological contribution by proposing a multiobjective optimization framework that explicitly linked irrigation decisions, groundwater extraction, and crop specific water requirements. Their work provided the theoretical basis for subsequent applications of multiobjective techniques in regulated irrigation systems and informed later case studies in different geographical contexts.
In the Bangladeshi context, Ullah and Nehring [26] applied a multiobjective optimization framework to the Muhuri Irrigation Project (MIP), using groundwater pumping limits and target environmental flow parameters adapted from the framework of Lewis and Randall. While their study demonstrated the feasibility of applying multiobjective methods to the MIP, it required substantial computational effort and explored only a limited portion of the feasible trade-off space.
The present study builds upon this line of research while introducing several methodological advancements. Unlike earlier studies that relied primarily on multiobjective formulations, this research independently analyzes two single objective models—maximizing economic benefit and minimizing environmental flow deficiency—under a common set of operational constraints. This approach enables the identification of extreme solutions that define the boundaries of the feasible solution space. Furthermore, alternative deterministic optimization techniques, including interior point and sequential quadratic programming methods, are employed to improve computational efficiency and reproducibility. While groundwater pumping limits and environmental flow targets are consistent with earlier studies, their applicability is justified by the hydrological, institutional, and operational characteristics of the Muhuri Irrigation Project, including seasonal water scarcity, centralized management, and the need to prevent saline intrusion in a coastal river system.
Given the limitations of existing models, particularly in the context of the Muhuri Irrigation Project, this study proposes efficient multiobjective optimization implemented techniques, including ‘gamultiobj’, ‘paretosearch’, and scalarization approaches. By integrating economic returns, crop specific water requirements, seasonal water availability, and explicit environmental flow constraints within a unified optimization framework, the study advances irrigation water allocation modeling and bridges the gap between theoretical formulations and practical, field level decision making in Bangladesh’s irrigation sector.
4 Problem Statement
Despite these objectives and its socioeconomic importance, the performance of the MIP has fallen short of initial expectations. During the dry winter months, surface water availability from the Muhuri River declines sharply, while rainfall contributes only about 2% of the annual total [26]. Although farmers tend to prefer Rabi crops because of their lower water requirements compared to Kharif crops, the limited availability of surface water during this period restricts the effective use of irrigation. As a coping strategy, many farmers increasingly depend on groundwater extraction, which raises production costs and places additional pressure on already stressed aquifers.
Institutional and infrastructural limitations further exacerbate these challenges. Inadequate water distribution systems, poor drainage conditions, and unplanned cropping intensities reduce the effectiveness of available irrigation water. Moreover, the centralized operation of the canal network, combined with limited control over on farm water delivery, often results in inequitable access to water, particularly disadvantaging farmers located at the tail ends of canals. Competition among users intensifies during dry periods, increasing the likelihood of inefficient and uneven water use. From an ecological perspective, reduced river inflows caused by irrigation withdrawals contribute to Environmental Flow Deficiency (EFD), threatening aquatic ecosystems and local fisheries that depend on sufficient freshwater availability. At the same time, excessive reliance on groundwater undermines long term sustainability by accelerating aquifer depletion and increasing energy consumption.
These interconnected socioeconomic, institutional, and environmental challenges highlight the limitations of existing operational practices within the Muhuri Irrigation Project. They underscore the need for an updated and integrated planning framework that explicitly accounts for both economic objectives and ecological constraints. A robust decision support system capable of guiding irrigation water allocation under conditions of seasonal scarcity is therefore essential to improve equity, provide a comprehensive representation of Environmental Flow Deficiency, support informed decision making by local authorities, enhance agricultural productivity, and ensure the long term sustainability of the Muhuri Irrigation Project.
5 Formulations of the Mathematical Models and Solution Approach
An irrigation planning authority must allocate land and water resources efficiently across multiple crops and time periods, ensuring economic viability and environmental sustainability. Surface water is delivered through canals, and additional water can be pumped. The system must meet staple crop area and production demands while minimizing environmental flow deficiency (EFD).
Decision variables:
$X_{c}$ Crop Area for each crop c in hectare haha
$Env.Flow_{m}$ Environmental Flow in month m in gigaliter GLGL
Input parameters:
$P_{c}$ Price of total crop production in $AUD$ per hectare (AUD/ha)
$C_{w}$ Cost of Surface Water conveyance (AUD/GL)
$K_{cm}$ Crop coefficient in the different growth stages of crops c in month m
$ET_{m}$ Evapotranspiration($GL/ha$) in each month m
$R_{m}$ Rainfall($GL/ha$) in each month m
$Allocation_{m}$ Surface Water($GL$) in each month m
$P_{m}$ Ground or Pumping Water($GL$) in each month m
$C_{p}$ Cost of pumping Groundwater(AUD/GL)
$Vcost_{c}$ Variable cost for each crop c per hectare (AUD/ha)
$Inflow_{m}$ Total runoff or Available River water ($GL$)in each month
$Tar.Env.Flow_{m}$ The minimum required Environmental flow($GL$) in each month
$T_{\text{pump}}$ Total Pumping Water $(GL)$
$T_{\text{area}}$ Total Area $(ha)$
$min_{area}$ Minimum Area $(ha)$
Decision variables dependent parameters:
$Allocation_{m}$ Allocated Surface water, where $Allocation_{m}={Inflow}_{m}-{Env.Flow}_{m}$
$P_{m}$ Pumped Ground Water, where $P_{m}=\sum_{c}(K_{cm}\cdot ET_{m}-R_{m})X_{c}-Allocation_{m}$
The objective functions can be formulated as follows:
The first objective, $f_{1}$:
$NB_{\text{max}}=\sum_{c}P_{c}X_{c}-C_{w}\cdot\sum_{m}\left(\sum_{c}(K_{cm}\cdot ET_{m}-R_{m})X_{c}-P_{m}\right)\\ -C_{p}\cdot\sum_{m}P_{m}-\sum_{c}Vcost_{c}\cdot X_{c}$ (1)
The second objective, $f_{2}$:
$EFD_{\text{min}}=\sum_{m}\max\left(\text{Tar.Env.Flow}_{m}-\text{Env.Flow}_{m},\ 0\right)$ (2)
Model 1: When Net Benefit is maximized
$\begin{array}[]{rl}\displaystyle\max_{(X_{c},Env.Flow_{m})\in X}&\ f_{1}(X_{c},Env.Flow_{m})\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (3)
where the feasible set $X$ is defined as
$\hskip-14.22636ptX:=\left\{(X_{c},Env.Flow_{m})\in\mathbb{R}^{c}\times\mathbb{R}^{m}\mid\sum_{m}P_{m}\leq T_{pump},\;\;min_{area}\leq X_{c},\;\;\sum_{c}X_{c}\leq T_{area}\right\}.$
The model seeks to maximize net benefit, with $X_{c}$ and $Env.Flow_{m}$ as the decision variables and parameters, including $P_{c}$, $C_{w}$, $K_{c,m}$, $ET_{m}$, $R_{m}$, $Allocation_{m}$, $P_{m}$, $C_{p}$, and $Vcost_{c}$. In this setup, maximum net benefit is achieved by sacrificing environmental performance, illustrating the trade-off between economic gain and ecological sustainability.
Model 2: When Environmental Flow Deficiency is minimized
$\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ f_{2}(X_{c},Env.Flow_{m})\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (4)
Model 2 is formulated as a single-objective optimization problem in which minimizing Environmental Flow Deficiency (EFD) is the sole objective. The target environmental flow $\text{Tar.Env.Flow}_{m}$ represents a minimum ecological requirement rather than an upper bound. Accordingly, the decision variable $\text{Env.Flow}_{m}$ may exceed this target when sufficient water is available or when environmental protection is prioritized. Any release above the target does not contribute to EFD, which is defined only as the shortfall below $\text{Tar.Env.Flow}_{m}$. Once the minimum environmental flow requirement is satisfied, EFD becomes zero for all feasible solutions.
For this reason, Model 2 is not intended to generate a realistic cropping plan in isolation. Instead, it serves as a boundary single-objective model that identifies the ecologically optimal extreme of the feasible solution space. In this formulation, crop allocation is governed by feasibility constraints such as minimum crop area requirements, total land availability, and pumping capacity rather than by economic incentives. The resulting solution is used as a reference point for evaluating trade-offs between economic and environmental objectives in the multiobjective formulation of Model 3.
Model 3: Multiobjective - Net Benefit versus Environmental Flow Deficiency
$\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\displaystyle\ \left\{-f_{1},f_{2}\right\}\\ \mbox{subject to the constraints}\\ &\displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}$ (5)
This multiobjective programming model simultaneously minimizes two conflicting goals: the negative Net Benefit function $f_{1}(X_{c},Env.Flow_{m})$ and the environmental flow deficiency function $f_{2}(X_{c},Env.Flow_{m})$. The constraints ensure feasible pumping limits, enforce both minimum and maximum land use requirements, and guarantee that crop area and environmental flow allocations remain non-negative. In essence, the formulation captures the trade-off between maximizing agricultural returns and preserving ecological water requirements.
We develop code within the MATLAB environment and utilise MATLAB solvers to address the models. A comprehensive solution procedure and algorithm for solving Models 1–3 are detailed in Appendices 10.1 and 10.2.
6 Numerical Experiments, Results and Sensitivity Analysis
This section presents a numerical study demonstrating that optimal irrigation water allocation can be achieved under a large number of operational constraints during dry year conditions. To account for the stochastic nature of the optimization solvers and their sensitivity to initial conditions, multiple randomized trials are performed. The solution procedure described in (10.1) is applied to solve the single objective formulations, where Model 1 (3), maximizes net economic benefit, and Model 2  (4), minimizes environmental flow deficiency. The single objective problems are solved using MATLAB nonlinear optimization solvers, including ‘fmincon’ implemented with the interior point method and the sequential quadratic programming (SQP) algorithm. For the multiobjective formulation, MATLAB’s genetic algorithm and Pareto search approach are employed, together with the scalarization method described in [3], to approximate the Pareto front and explore trade-offs between economic and environmental objectives. These solvers are evaluated based on their ability to handle large scale optimization problems and the challenges associated with approximating the Pareto front. All optimization experiments are conducted with randomly generated initial decision vectors and repeated multiple times to ensure robustness. For each solver and model combination, computational time and solution quality are recorded and summarized using representative statistics. All numerical computations are carried out on an HP Evo laptop equipped with $16$ GB RAM and a Core $i7$ processor operating at $4.6$ GHz.
Case Study 6.1
The numerical example is based on the framework developed by Ullah and Nehring [26], who applied an optimization model to the Muhuri Irrigation Project using dry-year data representative of low-flow, water-stressed conditions in Bangladesh. In the present study, dry-year scenarios are interpreted as operational worst-case conditions for irrigation planning, rather than as indicators of arid or climatically comparable environments. Accordingly, the modeling assumptions adopted here ensure methodological consistency and comparability with prior studies, while remaining aligned with the hydrological and institutional setting of the Muhuri Irrigation Project.
The analysis considers a regional water authority responsible for managing the ten most demand-intensive crops over a twelve-month planning horizon. The total cultivable area is 23,076 hectares, with a maximum annual groundwater pumping capacity of 50 GL and a minimum crop area requirement of 1,000 hectares per crop. The unit costs of pumped and surface water are set at 100,000 and 26,000 AUD per GL, respectively. The use of AUD, rather than Bangladeshi Taka (BDT) or USD, is maintained to ensure direct comparability with prior work. The target environmental flow is specified as 100 GL per month.(A fixed monthly environmental flow target is used to maintain consistency with the benchmark framework and to enable transparent comparison across alternative optimization formulations. Although environmental flow requirements in monsoon- dominated river systems are inherently seasonal, introducing seasonally varying targets would add hydrological complexity beyond the scope of this methodological study. Future work may incorporate seasonally adaptive environmental flow requirements.)
Key parameter values, including pumping limits and environmental flow targets, are adopted directly from existing studies [26, 27] to enable consistent benchmarking of alternative optimization formulations under identical operating conditions. These parameters are not intended as site-specific policy prescriptions, but rather as representative constraints for evaluating solution behavior and trade-off characteristics. All hydrological inputs, including river inflow, rainfall, and evapotranspiration, are adopted directly from the published datasets used in Ullah and Nehring [26], which were compiled from official project records of the Muhuri Irrigation Project. These data are treated as exogenous inputs to ensure methodological consistency and to focus the analysis on optimization behavior rather than hydrological estimation. Within this framework, allocation outcomes are examined both with and without enforcing the environmental flow constraint to identify the upper bound of attainable economic benefit and to explore trade-offs between profit maximization and environmental flow protection.
Table 1: Optimal Allocation of Crops and Environmental Flow for Model 1 (3).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 2076 1000 1000 5000 5000 5000
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 0 0 0 0 0 0 0 0 0 0 0 0
EFD 100 100 100 100 100 100 100 100 100 100 100 100
We solved Model 1 (3) and obtained the crop and environmental flow allocations shown in Table 1. In this case, the maximum profit achieved is $\mathdollar 1.4968\times 10^{9}$. However, since the second objective of maintaining environmental flows was fully relaxed, no water was allocated to environmental requirements, and no pumping was required. As a result, the environmental flow deficit reaches its maximum possible value of 1200 GL
Now we consider that the environment flow meets the target flow requirement of the environment in Model 2 (4), and Model  3 (5), that is,
$Tar.Env.Flow_{m}\leq Env.Flow_{m},$ (6)
Table 2: Optimal crop and environmental flow allocation for Model 2 (4) under constraint (6).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 2076 1000 1000 5000 5000 5000
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 100 100 100 100 100 100 100 100 100 100 100 100
EFD 0 0 0 0 0 0 0 0 0 0 0 0
We solved Model 2 (4) using constraint (6) and obtained the crop and environmental flow allocations presented in Table 2. In this scenario, the maximum profit is $\mathdollar 1.4080\times 10^{9}$. Here, the environmental flow deficiency is 0 GL, since constraint (6) does not allow for any shortfall. To satisfy this requirement, additional water must be supplied through pumping when surface water is insufficient.
The crop area allocations reported in Tables 1 and 2 exhibit relatively tidy values, with several crops assigned areas close to the minimum allowable level. This outcome is a direct consequence of the modeling structure rather than an artefact of the solution process. In particular, the presence of minimum crop area constraints, together with upper bounds on individual crop areas and the absence of any objective term promoting crop diversification, leads the optimization algorithms to select boundary solutions. Once feasibility conditions are satisfied, continuous nonlinear solvers such as ‘fmincon’ typically converge to corner points of the feasible region. As a result, the reported allocations should be interpreted as constraint-driven extreme solutions that define the bounds of the feasible space, rather than as detailed operational cropping plans.
Table 3: Optimal crop and environmental flow allocation for Model 3 (5).
crops Aus Aman Boro Wheat Potato Oil Pulses Suger Winter Summer
rice rice rice seeds cane vegi vegi
$X_{c}$ 1000 1000 1000 1000 1000 1000 1000 1000 1000 1000
Month Jan Feb Mar April May June July Aug Sep Oct Nov Dec
E.Flow 200 200 200 200 200 200 200 200 200 200 200 200
EFD 0 0 0 0 0 0 0 0 0 0 0 0
Next we minimize Model 3 (5) with the same parameter settings as in Case Study 6.1, considering both the inclusion and exclusion of constraint (6). The optimal crop and environmental flow allocations, shown in Table 3, are identical in both situations. In this case, the total environmental flow deficiency is zero, and the profit obtained is $\mathdollar 0.2742\times 10^{9}$. In Table 3, the environmental flow exceeds the target value of 100 GL per month and is uniformly allocated at 200 GL, reflecting surplus environmental releases under the multiobjective formulation rather than a change in the target environmental flow requirement. Note: The computational time taken to solve all three cases is approximately 8–10 seconds.
Observations from the Results: The three scenarios deliver different outcomes because each goal controls how water is shared between crops and the river:
  • Profit only: When the model focuses only on profit, nearly all available water is directed to the most valuable crops. This generates the highest income of $\mathdollar 1.4968\times 10^{9}$, but the river receives too little water, creating a total environmental flow shortage of $1200$ GL.
  • Profit with environmental target: When the model must also meet the river’s target flow, it reserves enough water for the environment. With less surface water left for irrigation, more pumping is needed, which is expensive. As a result, profit drops moderately to $\mathdollar 1.4080\times 10^{9}$, but the environmental shortage is reduced to zero.
  • Minimum environmental shortage: When protecting the environment is the top priority, the model sends a steady 200 GL each month to the river before allocating water to crops. Because of this and limits on pumping, not all land can be irrigated, and profit falls sharply to $\mathdollar 0.2742\times 10^{9}$. The river’s flow target, however, is fully met.
Overall, these results highlight a clear trade-off: directing more water to crops increases income but harms ecological health, while stronger environmental protection lowers profit.
7 Multiobjective Optimization Problems
Multiobjective optimization deals with finding the best balance between two or more objectives that often compete with each other. Because these goals conflict, there is rarely a single solution that is best for every objective at once. Instead, the aim is to identify a group of equally valid trade-off solutions called the Pareto front, where improving one objective would worsen at least one other. A common way to address such problems is scalarization, which reformulates a multiobjective problem as a single-objective one so that standard optimization techniques can be applied. In this work, we use the weighted-constraint method [3]. Here, the objective functions are expressed with adjustable weights: one weighted objective is optimized, while the remaining objectives are treated as constraints whose bounds depend on the chosen weights. By varying these weights evenly, we can approximate the complete Pareto optimal set. We chose the weighted constraint approach for its straightforward implementation and its ability to generate accurate Pareto solutions (more scalarization approaches in [4], [22]). The detailed procedure for the two objective case is described below.
The weighted-constraint approach ($P_{w}^{k}$) : For a positive weights we define
$\hskip 56.9055ptW:=\left\{w\in\mathbb{R}^{2}\mid w_{1},w_{2}>0,\sum_{i=1}^{2}w_{i}=1\right\}.$
Let $w\in W$, $f_{1}$ in (1), and $f_{2}$ in (2), and we therefore define two scalar problems as
$\mbox{($P_{w}^{1}$)}\ \left\{\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ \ w_{1}f_{1}(X_{c},Env.Flow_{m}),\\[11.38109pt] \mbox{subject to}&\ \ w_{2}f_{2}(X_{c},Env.Flow_{m})\leq w_{1}f_{1}(X_{c},Env.Flow_{m}),\\ &\ \ \displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}\right$. (7)
and
$\mbox{($P_{w}^{2}$)}\ \left\{\begin{array}[]{rl}\displaystyle\min_{(X_{c},Env.Flow_{m})\in X}&\ \ w_{2}f_{2}(X_{c},Env.Flow_{m}),\\[11.38109pt] \mbox{subject to}&\ \ w_{1}f_{1}(X_{c},Env.Flow_{m})\leq w_{2}f_{2}(X_{c},Env.Flow_{m}),\\ &\ \ \displaystyle X_{c},Env.Flow_{m}\geq 0.\end{array}\right$. (8)
In Section 6, net profit and environmental deficiency are optimized separately, and the results highlight the trade-off between these two objectives. In Case Study 7.1, we solve the two objective optimisation problem defined in (5).
Case Study 7.1
In this analysis, we use the input parameters from Case Study 6.1 to solve the multiobjective optimization problem defined in Model 3. Because it is a large scale problem, different approaches are applied, including the ’gamultiobj’ and ’paretosearch’ solvers in the MATLAB environment, while the weighted-constraint method is employed to approximate the Pareto front.
We solve the problem described in (5) using the ‘gamultiobj’ solver, increasing the population size up to $500$. The lower bounds are set as the minimum area for $X_{c}$ and zero for environmental flow, while the upper bounds are set at $5,000$ ha for $X_{c}$, (The per-crop upper bound is implemented as it ensures realistic crop distribution while preventing extreme monocropping and preserving comparability with the benchmark framework) and $300$ GL for environmental flow. The Pareto front obtained with this algorithm is shown in Figure 1(a), and the computation time along with the total number of Pareto points is reported in Table 4.
We now solve the problem defined in (5) using the ‘paretosearch’ solver, with the population size increased to 500. The same lower and upper bounds are considered for $X_{c}$ and environmental flow values. The Pareto front approximated by this algorithm is shown in Figure 1(b), while the computation time and the total number of Pareto points are reported in Table 4.
Finally, we use the weighted constraint scalarization approach. This method requires a sequence of steps detailed in Appendix 10.2. With weight normalisation as defined in $W$, the procedure involves solving subproblems (7) and (8) and comparing their solutions, discarding any non-dominated points. To solve the subproblems, we employ the fmincon solver with the ‘Interior Point’, ‘SQP’, and ‘SQP Legacy’ algorithms, using the same lower and upper bounds as in the previous two cases. Following the algorithm outlined in Appendix 10.2, we then solve the problem defined in (5) to approximate the Pareto points, which are illustrated in Figure 1(c). A numerical summary is provided in Table 4.
Table 4: Numerical performance of ’gamultiobj’, ’paretosearch’ and $P_{w}^{k}$.
Methods Number of Population Number of Time Pareto Non-Pareto
grid points size subproblems in seconds points points
gamultiobj 500 15.72 345 Nil
paretosearch 500 16.86 500 Nil
$P_{w}^{k}$ 500 1000 10.12 989 11
(a) Pareto front approximated with ’gamultiobj’.
(b) Pareto front approximated with ’paretosearch’.
(c) Pareto front approximated with ($P_{w}^{k}$).
Figure 1: Pareto fronts generated by the ’gamultiobj’, ’paretosearch’, and ’weighted-constraint’ ($P_{w}^{k}$) methods.
Comparison Analysis: We observe that the ‘gamultiobj’ algorithm generates $345$ Pareto points in $15.72$ seconds under the same parameter settings. From Figure 1(a), the highest environmental deficiency level is around $130$ GL when profit is $\mathdollar 1.91\times 10^{9}$, and the deficiency drops to zero when profit is between ($\mathdollar 1.180\;\mbox{and}\;1.181)\times 10^{9}$. The ‘paretosearch’ algorithm takes slightly more time but produces more Pareto points $500$ in total. In generating the Pareto front, what matters more than the number of points is how much of the Pareto front each algorithm can cover. For paretosearch, the highest deficiency is $1200$ GL at a profit of $\mathdollar 1.49\times 10^{9}$, with deficiency reaching zero when profit is around $\mathdollar 1.37\times 10^{9}$ (see, Figure 1(b)). Using the weighted-constraint scalarization approach, we find that the highest deficiency and corresponding profit are the same as ’paretosearch’. However, the deficiency starts to appear when the profit is $\mathdollar 1.4\times 10^{9}$ and decreases to $\mathdollar 0.2\times 10^{9}$ (see, Figure 1(c)).
To evaluate the true Pareto front, we refer to Case Study 6.1 results, where individual objectives $f_{1}$ and $f_{2}$ are optimised. The results show that the highest deficiency of $1200$ GL occurs at a profit of $\mathdollar 1.49\times 10^{9}$, consistent with Figures 1(b) and (c) obtained for ‘paretosearch’ and the ‘scalarization’ approach, respectively. Moreover, maximising $f_{1}$ under target environmental flow constraints (6) in Case Study 5.1 results in zero deficiency when profit is $\mathdollar 1.4\times 10^{9}$, which matches only the scalarization results in Figures 1(c). This indicates that ‘paretosearch’ and ‘gamultiobj’ fail to approximate the lower part of the Pareto front. Therefore, considering computation time, the number of generated points, and the accuracy of the true Pareto front, the scalarization approach performs better than the other methods.
Comparison analysis with the previous study[26]: The findings of the present study highlight several critical improvements over the earlier work of Ullah and Nehring [26]. A key point of divergence lies in the treatment of Environmental Flow Deficiency (EFD). Ullah and Nehring reported a maximum EFD of only 35 GL under a fixed monthly environmental flow target of $100$ GL. Given that EFD is defined as the shortfall between the prescribed environmental flow and the actual allocation, the theoretical maximum deficit for a uniform 100 GL monthly target is $1200$ GL per year (see, obtained results in Table 1). This indicates that their model captured only a narrow portion of the feasible solution space. Consequently, the ecological shortfalls are understated, and the trade-offs between agricultural benefits and ecological requirements are not fully explored. In contrast, our scalarization based approach succeeds in covering the entire feasible frontier, generating Pareto solutions that extend up to the theoretical maximum EFD. This broader coverage, as illustrated in Figure 1(c) (Pareto fronts generated by the Weighted-Constraint approach, $P_{w}^{k}$), provides a far more comprehensive representation of the trade-off surface compared to [26, Figure 1 in Page 10], where the solution set remains confined to a very limited region.For cross method comparison, relative performance is discussed based on the shape and extent of the attainment surfaces rather than absolute magnitudes alone. The two studies also differ markedly in terms of computational performance and solution diversity. Ullah and Nehring, using the NSGA-II evolutionary algorithm, reported only $34$ Pareto optimal solutions, obtained at the expense of approximately $706.59$ minutes of computational time. In sharp contrast, the scalarization technique produces $989$ Pareto solutions in just $10.12$ seconds. This demonstrates not only substantial computational efficiency but also a significant improvement in the diversity of solutions available to decision makers. The larger and more representative solution set provided by our approach ensures a more reliable understanding of trade-offs and expands the range of strategies that can be considered in irrigation planning.
Together, these results establish that the scalarization framework achieves superior performance in both accuracy and efficiency. By covering the full spectrum of environmental deficits and delivering a wide set of optimal solutions with minimal computational burden, our study offers a more rigorous and practical foundation for decision making in the management of the Muhuri Irrigation Project.
8 Conclusion
This study contributes to the growing efforts in sustainable water management by introducing a practical optimization framework designed to strike a balance between economic gain and environmental responsibility. By applying three widely used multiobjective optimization approaches, implemented through the MATLAB functions ‘gamultiobj’ and ‘paretosearch’ alongside a scalarization method, the framework provides a transparent means of exploring trade-offs in irrigation water allocation under scarcity. The findings suggest that smart allocation strategies can lead to noticeably higher economic benefits while also safeguarding environmental flows. The results show that scalarization, in particular, is well suited to handle trade-offs between conflicting goals. The model can also successfully consider real-world operational constraints, making it relevant and applicable in practice. Moreover, since the analysis is focused on a single irrigation project in a tropical monsoon area, the findings are also applicable in arid and semi arid areas. The primary contribution of this study lies in methodological comparison and optimization performance assessment rather than direct operational prescription.
The study has several limitations related to data and modeling assumptions. It adopts a deterministic framework in which input data are assumed to be known with certainty and relies solely on dry year conditions. As a result, the model does not capture uncertainties associated with climate variability, crop yield fluctuations, or changes in target environmental flow requirements. Moreover, fuzzy and socio-policy dynamics are not explicitly represented.
There is substantial scope to further enhance the proposed modeling framework by incorporating additional operational and hydrological complexities. Future work may include crop specific area constraints based on demand, canal capacity limits, and crop scheduling decisions linked to land suitability and crop types. The model can also be extended to represent key hydrological processes such as deep percolation, soil moisture dynamics, dam storage, overflow, and reservoir evaporation, particularly during periods of excess rainfall. Introducing these elements would improve the representation of water availability and storage behavior under monsoon conditions. In addition, allowing the target Environmental Flow to vary seasonally would better reflect natural hydrological variability and improve ecological realism. Such seasonally adaptive targets would enable a more accurate assessment of trade-offs between irrigation withdrawals and downstream environmental protection across wet and dry periods. Finally, integrating the framework with socioeconomic and policy modules would enhance its relevance for decision makers at both local and national levels.
Acknowledgment: We acknowledge ChatGPT (5) AI tools, which are used to rephrase, edit, and polish the author’s written text for spelling, grammar, or general style.
9 Funding and/or Conflicts of interests/Competing interests
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript. The authors have no relevant financial or non-financial interests to disclose.